Manuel is a college student who lives in Houston and provides math helper for extra cash

11. Calculating the price elasticity of supply


Manuel is a college student who lives in Houston and provides math helper for extra cash. At a wage of $50 per hour, he is willing to help 7 hours per week. At $65 per hour, he is willing to tutor 10 hours per week.


Using the midpoint method, the elasticity of Manuel’s labor supply between the wages of $50 and $65 per hour is approximately   , which means that Manuel’s supply of labor over this wage range is   .
Answer #1

Using the mid-point method, the calculation of elasticity is given by Ed = {(Q2-Q1)/Q}/{(P2-P1)/P}, where Q = (Q1 + Q2)/2 and P = (P1 + P2)/2

Here, P = (50+65)/2 and (P2-P1)/P = 15*2/135

Q = (7+10)/2 and (Q2-Q1)/Q = 3*2/17

Ed = 6/17 * 135/30 = 1.58

The elasticity is approximately 1.58 which means that Manuel’s supply of labor over this wage range is relatively elastic in nature.